Suppose that the continuous random variable x has the probability density function


f(x) = 1/2e−|x| , −∞ < x < ∞


find the m.g.f of x and use it to find E(x) and V ar(x).

1
Expert's answer
2022-04-27T14:27:43-0400
"M(t)=\displaystyle\int_{-\infin}^{\infin}e^{tx}f(x)dx"

"=\dfrac{1}{2}\displaystyle\int_{-\infin}^{0}e^{tx}e^xdx+\dfrac{1}{2}\displaystyle\int_{0}^{\infin}e^{tx}e^{-x}dx"

"=\dfrac{1}{2(t+1)}[e^{(t+1)x}]\begin{smallmatrix}
 0 \\
 -\infin
\end{smallmatrix}+\dfrac{1}{2(t-1)}[e^{(t-1)x}]\begin{smallmatrix}
 \infin \\
 0
\end{smallmatrix}"

"=\dfrac{1}{2(t+1)}(1-0)+\dfrac{1}{2(t-1)}(0-1)"

"=\dfrac{1}{1^2-t^2}"

"M(t)=\dfrac{1}{1^2-t^2}"

Differentiate MGF with respect to t.


"M'(t)=\dfrac{2t}{(1^2-t^2)^2}"

Put "t=0"


"E(X)=M'(0)=0"

Find "M''(t)"


"M''(t)=2\cdot\dfrac{(1^2-t^2)^2-2t(1^2-t^2)(-2t)}{(1^2-t^2)^4}"

"=2\cdot\dfrac{1^2+3t^2}{(1^2-t^2)^3}"

Put "t=0"


"E(X^2)=M''(0)=2"


"Var(X)=E(X^2)-(E(X))^2"

"=2-(0)^2=2"

"M(t)=\dfrac{1}{1^2-t^2}"


"E(X)=0"


"Var(X)=2"